OpenStudy (jordan123321):

What is the value of x? x over 4 equals 14 over 16 x =

OpenStudy (jordan123321):

help dont just sit there

OpenStudy (k_lynn):

@jordan123321 do you still need some help?

OpenStudy (jordan123321):

yes

OpenStudy (k_lynn):

Ok. The question says Over two times. That means that we're going to set up fractions. Or a proportion. So x Over 4 would be \[\frac{ x }{ 4 }\] and 14 Over 16 would be \[\frac{ 14 }{16 }\] Right?

OpenStudy (jordan123321):

yes

OpenStudy (k_lynn):

So now that we have the fractions, We just need an = sign in between and solve for x. \(\Large\frac{ x }{ 4 }\) = \(\Large\frac{ 14 }{16 }\) We can solve for x by cross multiplying and then dividing by what's left. To do that, multiply the numbers that are across from each other and Then multiply by the number you have left. I don't feel like that's very helpful, so I'll show you what I mean.